OpenStudy (anonymous):

John has a total of 13 coins in his pocket. Of the 13 coins, they are all either quarters or dimes. The total value of the coins is $2.50. How many quarters does John have in his pocket?

Parth (parthkohli):

If john has x dimes, then he has 13 - x quarters(because x + 13 - x = 13)

Parth (parthkohli):

Let's make an equation here. \( \color{Black}{\Rightarrow x(10) + (13 - x)(25) = 250 }\) I expressed the $2.50 figure as cents.

OpenStudy (anonymous):

So, how would I figure that?

OpenStudy (anonymous):

The first step in solving this problem is to build two equations. 0.25Q+0.10D=2.50 and Q+D=13 where Q=total number of quarters and D=total number of Dimes. Next we try to subtract one equation from the other to eliminate one variable. So we multiply 0.25Q+0.10D=2.50 by 4 to get 1Q+0.4D=10. Then we subtract Q+0.4D=10 from Q+D=13. This results in 0.6D=3. We divide both sides by 0.6 and end up with d=5. Therefore we have 5 dimes, which equals 50 cents. Therefore we have $2 in quarters or 8 quarters. 8 quarters is the answer.

OpenStudy (anonymous):

good enough

OpenStudy (anonymous):

did you really understand it ??? can you solve other such questions ???

OpenStudy (anonymous):

Sort of

OpenStudy (anonymous):

ok if i give you a similar question can you solve it ???

OpenStudy (anonymous):

Maybe

OpenStudy (anonymous):

I can try

OpenStudy (anonymous):

ok so here is a question:

OpenStudy (anonymous):

In a culture plate the number of bacteria initially was 2 times the number of virus in a second plate and there total sum is 50000. after an hour the their to total number becomes 100000 and in culture plate one bacteria become 4 times but the virus remains the same. calculate the original no of bacteria and virus ?

OpenStudy (anonymous):

@Kittyrae15 can you solve it

OpenStudy (anonymous):

ok let me make you the eq no of virus = x no of bacteria = y x+2y=50000 after an hour total = 100000 4(2y)+x=100000 x+8y=100000 now by substitution or elimination methods solve it.